OpenStudy (nermaljean99):

Help on Algebra 2???!!! :) <3

OpenStudy (nermaljean99):

10.) The length of the sides of a rectangular window have the ratio 1.7 to 1. The area of the windows is 2,720 square inches. What are the dimensions of the window? (1 point) a.) 40 inches by 36 inches b.) 40 inches by 65 inches c.) 40 inches by 34 inches d.) 40 inches by 68 inches

OpenStudy (nermaljean99):

@dan815

OpenStudy (alyssa_xo):

draw out the rectangle, then write out the lengths in terms of x.

OpenStudy (alyssa_xo):

that's a dumb way of going about things. you should be able to find the lengths without looking at the choices

OpenStudy (alyssa_xo):

rude

OpenStudy (nermaljean99):

@tocuteforyou don't you have better things to than be little people? seriously. smh some people need to get a life

OpenStudy (nermaljean99):

besides this isn't a website based on how people look, its a homework help site!

OpenStudy (alyssa_xo):

it doesn't hurt to look pretty. I met my ex wife here

OpenStudy (nermaljean99):

So the answer is D? orrr what? XD i got a whole test to do

OpenStudy (alyssa_xo):

It's alright. What I did was \(1.7x^2=2720\) divide both sides by 1.7 \(x^2=1600\) square root both sides x=40 1.7x=40*1.7=68

OpenStudy (nermaljean99):

and boom goes the dynamite XD lol thanks

OpenStudy (nermaljean99):

oh and @tocuteforyou I have been told many times im pretty... so ya

OpenStudy (anonymous):

your very pretty @nermaljean99

OpenStudy (nermaljean99):

Thanks!

OpenStudy (alyssa_xo):

marry me @nermaljean99

OpenStudy (nermaljean99):

lol XD

OpenStudy (nermaljean99):

I have a few more questions

OpenStudy (alyssa_xo):

close this question, post another, tag me, marry me

OpenStudy (nermaljean99):

lol ok